# Insulator phases of Bose–Fermi mixtures induced by intraspecies next-neighbor interactions

### v2

@@ 3 Fermionic next-neighbor interactions
The key findings of this investigation are shown in the phase diagrams from Fig. 1(a), where the different insulator phases that can appear at a given filling are summarized. Each diagram assumes that all densities attain rational values, which guarantees the commensurability between the chain and each component density. This schematic includes the results already known for the single-component models: for $\rho_{\rm B}=0$, we have the band insulators at $\rho_{\rm F}=0$ and $2$ and the Mott insulator at half-filling $(\rho_{\rm F}=1)$, well known from the original Fermi-Hubbard model[88]. On the other hand, for $\rho_{\rm F}=0$, we find the analogous band insulator at $\rho_{\rm B}=0$ and the bosonic Mott insulator at unit-filling $(\rho_{\rm B}=1)$ [89, 90, 91, 70, 92]. In the case of $V_{\rm BB}\neq 0$ (Fig. 1(b)) the CDW insulator is found for bosonic chains with next-neighbor interactions at $\rho_{\rm B}=1/2$, which has been found in previous papers [72, 48, 73].
Now we focus our analysis on nonzero fermionic next-neighbor interactions only at three essential fermionic densities, given by $\rho_{\rm{F}}=2/3$, $2/5$, and $4/5$ at zero population imbalance $I=0$, showing for each of them the emergence of an unknown insulator, along with a discussion on how the next-neighbor interactions between fermions affect the already well-known insulator phases. For these investigations, we fix the on-site interactions at $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$, while we choose a next-neighbor interaction that will be within the range $0\leq V_{\rm{FF}}\leq 6$. It has been shown in previous studies that the presence of a next-neighbor interaction in the model leads to finite-size effects that perturb the ground-state energy if the number of sites is even [70]; hence we restrict $L$ to always being odd. We also look at the imbalanced case $I\neq 0$ and how this affects the insulating phases of the model.
@@
3.1 Two-thirds fermionic filling $(\rho_{\rm{F}}=2/3)$
@@
Figure 2: (a) Bosonic density $\rho_{\rm{B}}$ vs. bosonic chemical potential $\mu_{\rm{B}}$ at the thermodynamic limit for a balanced mixture with a fermionic density of $\rho_{\rm{F}}=2/3$ and local interactions $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$. Here we compare the behavior with and without next-neighbor interaction between fermions, using $V_{\rm{FF}}=0$, $2$. The inset shows the bosonic charge gap $\Delta_{\rm{B}}$ as a function of the system size $L$ for each plateau found for $V_{\rm{FF}}=2$. In (b), we show the density profile for bosons $\langle\hat{n}^{\rm{B}}_{j}\rangle$ (red circles) and fermions $\langle\hat{n}^{\rm{F}}_{j}\rangle$ (blue diamonds) for $L=75$ at the plateau with $\rho_{\rm{B}}=1/2$ found in (a). The points correspond to DMRG results, and the lines are visual guides.
@@
For the turned-off next-neighbor interaction, we found a monotonous increase in $\mu_{\rm{B}}$ as we continuously added bosons to the system in the majority of the graph, which corresponds to the presence of superfluid (SF) phases, since they have nonzero compressibility. In particular, this behavior has exceptions at $\rho_{\rm{B}}=1/3$, $2/3$, where the graph is discontinuous, and we found incompressible phases. Each plateau in the $\rho_{\rm{B}}-\mu_{\rm{B}}$ curve represents the presence of an insulator state, where we denote its charge gap $\Delta_{\rm{B}}$ as the length of the plateau according to (6). At $\rho_{\rm{B}}=1/3$, we obtained a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=1.16$, corresponding to the well-known mixed Mott insulator (MMI), in which the interplay between the commensurability of both bosons and fermions and their mutual interaction generates the presence of an incompressible phase. On the other hand, for $\rho_{\rm{B}}=2/3$, a plateau with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.01$ appears, which we identify as a spin-selective Mott insulator (SSMI) that manifests an analog behavior, as does the previous phase, but in this one, the association is between bosons and only one type of fermion (spin up or down). These insulator phases are already known in the literature [64, 31, 65, 66, 67, 68] and follow the relations $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ for the MMI and $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ for the SSMI in a balanced Bose–Fermi mixture.
@@
After turning on the next-neighbor interaction, we found ourselves in a similar context, where most of the graph is in a SF phase. We still observed both previously-mentioned insulators, but now their charge gaps have changed. For the MMI, we found that the gap decreases to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=0.34$, which is less than a third of its original value. Homogeneous distribution of carriers is expected in the absence of next-neighbor interaction, making it expensive to increase the number of particles, but allowing non-local coupling between fermions leads to the formation of composites between bosons and fermions, which have a less homogeneous distribution of charge along the lattice, thus diminishing the energy gap. On the other hand, the SSMI slightly increases its gap to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.81$, showing that this phase benefits from the non-local fermion interaction. In this case, only half of the fermions are coupled to the bosons $\left(\rho^{\uparrow,(\downarrow)}_{\rm{F}}=1/3\right)$; however, the excess in the number of bosonic carriers makes the distribution of charge more homogeneous, making the effect of the next-neighbor interactions less dramatic.
@@
On observing Fig. 2(a) for $V_{\rm{FF}}=2$, we notice that an additional incompressible phase at $\rho_{\rm{B}}=1/2$ emerges with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=0.71$. To characterize this mysterious insulator, in Fig. 2(b) we show the corresponding bosonic and fermionic density profiles. Since the fermions repel each other, the optimal configuration is to leave space between them, which produces the characteristic oscillating pattern from the CDW phase. Moreover, the bosons also interact with the fermions locally, and hence they locate themselves between them, which generates another CDW order. Since the patterns are out of phase with each other, we call this insulator an immiscible CDW (ICDW). This curious insulator always appears at $\rho_{\rm{B}}=1/2$ (see Fig. 1(a)), since only for this bosonic density is the CDW structure stable for both species. That is, if there were more bosons, they could not fit between the fermions, and if there were fewer bosons, the fermions could flow freely through the spaces left behind by the missing bosons, removing the insulator condition. This shows that the ICDW represents the first example of an incommensurable insulator in the Bose–Fermi mixture model, a product of non-local interactions. Moreover, a similar bidimensional pattern was observed with neutral and charged dipolar excitons of $\rm{GaAs}$ bilayers recently[34], which allows us to confirm the relevance of the next-neighbor interactions in this Bose–Fermi system.
@@
Figure 3: (a) Phase diagram of the bosonic chemical potential $\mu_{\rm{B}}$ vs. next-neighbor fermion interaction magnitude $V_{\rm{FF}}$ at the thermodynamic limit for a fermionic density $\rho_{\rm{F}}=2/3$ without spin imbalance and local interactions $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$. Here, we show the evolution of the respective charge gaps from each insulator phase shown in Fig. 2(a) which include the MMI at $\rho_{\rm{B}}=1/3$, the SSMI at $\rho_{\rm{B}}=2/3$ and the ICDW at $\rho_{\rm{B}}=1/2$. The points are extrapolations of the chemical potential to the thermodynamic limit. (b) Order parameter of the ICDW phase $\mathcal{O}^{\rm{ICDW}}$ as a function of $V_{\rm FF}$ at the thermodynamic limit for the same parameters as (a). The vertical dashed line denotes the critical value ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/2}$ of the SF-ICDW transition obtained from the order parameter calculation. The inset shows the dependence of $\mathcal{O}^{\rm{ICDW}}$ on $\rho_{\rm B}$ for a chain of $L=45$ and for $V_{\rm FF}=0$ (black squares) and $V_{\rm FF}=2$ (red circles). Here, $\rho_{\rm B}=1/2$ is depicted as a dotted line. Horizontal dashed and dotted lines are added to highlight the zero value of the order parameter. The solid lines are visual guides.
@@
This signals the simultaneous CDW order of both species as the multiplication of the intraspecies CDW order parameters, with a minus sign representing the out-of-phase mutual oscillation, exhibited in Fig. 2(b). Moreover, we subtract the total density from each corresponding density profile to remove the contributions of flat profiles. We show $\mathcal{O}^{\rm{ICDW}}$ for $0\leq V_{\rm FF}\leq 2$ extrapolated to the thermodynamic limit in Fig. 3(b), where we can see the sudden change from zero to positive values after the critical point ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/2}\approx 0.7$ (indicated by the vertical line) which agrees with the one obtained from the gap calculation. The dependence on the number of bosons is also shown in the inset of Fig. 3(b) as we vary $\rho_{\rm B}$ and observe that the main nonzero contribution occurs close to half-filling for $V_{\rm FF}=2$ (red dots), while for $V_{\rm FF}=0$ (black squares), the order parameter is close to zero, as expected. Since we restrict ourselves to odd values of $L$, we observe a deviation of the maximum from $\rho_{\rm B}=1/2$; nevertheless, by increasing the system size this finite-size effect disappears as the maximum gets closer to the expected value.
@@
On the other hand, for the MMI at $\rho_{\rm{B}}=1/3$, we see that the charge gap decreases monotonously until it vanishes at a critical value of ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/3}\approx 3.5$, as anticipated from the formation of composites that destroy the crystal order mentioned beforehand. For the SSMI, even though we observed an initial increase in its charge gap, here we find that after a certain value, the gap starts to decrease monotonously. This behavior is persistent until the charge gap vanishes at a very high $V_{\rm{FF}}$ value, not shown here. Although the excess of bosonic carriers and the fact that they tie only one kind of fermion together try to maintain a homogeneous distribution of charge in the system, large values of the next-neighbor interaction will force the emergence of compounds that diminish the ground-state energy and finally destroy the crystalline order.
@@
It is worth noting that the ICDW phase does not appear for every fermionic density. Specifically, for $\rho_{\rm{F}}<1/2$, the system does not have enough fermions to build the sub-lattice for the CDW order. This is independent of the spin-population imbalance, since the CDW order can be established independently of the number of fermions in each orientation. On the other hand, for $I\neq 0$ there is another limitation. Consider the case where $\rho_{\rm{F}}^{\uparrow}$ or $\rho_{\rm{F}}^{\downarrow}$ are larger than half filling; then all of the fermions with the corresponding spin do not fit in the CDW order, and hence it is not possible to construct the insulating phase. Both restrictions can be summarized with the mathematical condition that the ICDW phase only appears for $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$. The case $I=1$, that is of spin-polarized fermions, has already been reported in the literature [83], and according to this study, there is only one possible density combination in which the insulator can appear, namely $\rho_{\rm{F}}=\rho_{\rm{B}}=1/2$. This emphasizes that a mixture with a larger number of fermionic degrees of freedom admits a more general family of density combinations for the emergence of the ICDW state.

### v3

@@ 3 Fermionic next-neighbor interactions
The key findings of this work are shown in the phase diagrams from Fig. 1, where the different insulator phases that can appear at a given filling are summarized. Each diagram assumes that all densities attain rational values, which guarantees the commensurability between the chain and each component density. Moreover, since all density relations that define the possible configurations a certain insulator can have are linear functions of $\rho_{\rm B}$, $\rho_{\rm F}^{\uparrow}$, and $\rho_{\rm F}^{\downarrow}$ (see Sec. 1), each insulator is represented by a straight line in the phase diagram. This schematic also includes the results already known for the single component models: For $\rho_{\rm B}=0$ we have the band insulators at $\rho_{\rm F}=0$ and $2$, and the Mott insulator at half-filling $(\rho_{\rm F}=1)$, well known from the original Fermi-Hubbard model[92]. On the other hand, for $\rho_{\rm F}=0$ we find the analogous band insulator at $\rho_{\rm B}=0$ and the bosonic Mott insulator at unit-filling $(\rho_{\rm B}=1)$ [93, 94, 95, 74, 96]. In the case of $V_{\rm BB}\neq 0$ (Fig. 1(b)) the CDW insulator is found for bosonic chains with next-neighbor interactions at $\rho_{\rm B}=1/2$, which has been found in previous works [76, 50, 77].
Now we focus our analysis on nonzero fermionic next-neighbor interactions only at three essential fermionic densities given by $\rho_{\rm{F}}=2/3$, $2/5$, $4/5$ at zero population imbalance $I=0$, showing for each of them the emergence of an unknown insulator, along with a discussion on how the next-neighbor interactions between fermions affect the already well-known insulator phases. For these investigations, we fix the on-site interactions at $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$, while we choose a next-neighbor interaction, which will be within the range $0\leq V_{\rm{FF}}\leq 6$. It has been shown in previous work that the presence of a next-neighbor interaction in the model leads to finite-size effects that perturb the ground-state energy if the number of sites is even [74], hence we restrict $L$ to be always odd. We also look at the imbalanced case $I\neq 0$ and how this affects the insulating phases of the model.
@@
3.1 Two thirds fermionic filling $(\rho_{\rm{F}}=2/3)$
@@
Figure 2: (a) Bosonic density $\rho_{\rm{B}}$ vs. bosonic chemical potential $\mu_{\rm{B}}$ at the thermodynamic limit for a balanced mixture with a fermionic density $\rho_{\rm{F}}=2/3$ and local interactions $U_{\rm{FF}}=4$, $U_{\rm{BF}}=6$. Here we compare the behavior with and without next-neighbor interaction between fermions, using $V_{\rm{FF}}=0$, $2$. The inset shows the bosonic charge gap $\Delta_{\rm{B}}$ as a function of the system size $L$ for each plateau found for $V_{\rm{FF}}=2$. In (b), we show the density profile for bosons $\langle\hat{n}^{\rm{B}}_{j}\rangle$ (red circles) and fermions $\langle\hat{n}^{\rm{F}}_{j}\rangle$ (blue diamonds) for $L=75$ at the plateau with $\rho_{\rm{B}}=1/2$ found in (a). The points correspond to DMRG results, and the lines are visual guides.
@@
For the turned-off next-neighbor interaction, we found a monotonous increase in $\mu_{\rm{B}}$ as we continuously add bosons to the system in the majority of the graph, which corresponds to the presence of superfluid (SF) phases since they have nonzero compressibility. In particular, this behavior finds its exceptions at $\rho_{\rm{B}}=1/3$, $2/3$ where the graph is discontinuous; we found incompressible phases. Each plateau in the $\rho_{\rm{B}}-\mu_{\rm{B}}$ curve represents the presence of an insulator state, where we denote its charge gap $\Delta_{\rm{B}}$ as the length of the plateau according to (8). At $\rho_{\rm{B}}=1/3$, we obtained a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=1.16$ corresponding to the well-known mixed Mott insulator (MMI) in which the interplay between commensurability of both bosons and fermions and their mutual interaction generate the presence of an incompressible phase. On the other hand, for $\rho_{\rm{B}}=2/3$, a plateau with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.01$ appears, which we identify as a spin-selective Mott insulator (SSMI) that has an analog behavior as does the previous phase, but instead, the association is between bosons and only one type of fermion (spin up or down). These insulator phases are already known in the literature [66, 31, 67, 68, 69, 70] and follow the relations $\rho_{\rm{B}}+\rho_{\rm{F}}=1$ for the MMI and $\rho_{\rm{B}}+\rho_{\rm{F}}/2=1$ for the SSMI in a balanced Bose–Fermi mixture.
@@
After turning on the next-neighbor interaction, we found ourselves in a similar context, where most of the graph is in a SF phase. We still observe both previously mentioned insulators, but their charge gaps have changed. For the MMI, we found that the gap decreases to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/3}=0.34$, less than a third of its original value. A homogeneous distribution of carriers is expected in the absence of next-neighbor interaction, making it expensive to increase the number of particles, but allowing long-ranged coupling between fermions leads to the formation of composites between bosons and fermions, which have a less homogeneous distribution of charge along the lattice, diminishing the energy gap. On the other hand, the SSMI slightly increases its gap to $\Delta_{\rm{B}}^{\rho_{\rm{B}}=2/3}=2.81$, showing that this phase benefits from the long-ranged fermion interaction. In this case, only half of the fermions are coupled to the bosons $\left(\rho^{\uparrow,(\downarrow)}_{\rm{F}}=1/3\right)$; however, the excess in the number of bosonic carriers makes the distribution of charge more homogeneous, making the effect of the next-neighbor interactions less dramatic.
@@
On observing Fig. 2(a) for $V_{\rm{FF}}=2$, we notice that an additional incompressible phase at $\rho_{\rm{B}}=1/2$ emerges with a charge gap of $\Delta_{\rm{B}}^{\rho_{\rm{B}}=1/2}=0.71$. To characterize this mysterious insulator, we show the corresponding bosonic and fermionic density profiles in Fig. 2(b). Since the fermions repel each other, the optimal configuration is to leave space between them, which produces the characteristic oscillating pattern from the CDW phase. Moreover, the bosons also interact with the fermions locally, hence they locate themselves between them, which generates another CDW order. Since the patterns are out of phase with each other, we call this insulator an immiscible CDW (ICDW). This curious insulator always appears at $\rho_{\rm{B}}=1/2$ (see Fig. 1(a)), since only for this bosonic density is the CDW structure stable for both species. That is, if there were more bosons, they could not fit between the fermions, and if there were fewer bosons, the fermions could flow freely through the spaces left behind by the missing bosons, removing the insulator condition. A similar bi-dimensional pattern was observed with neutral and charged dipolar excitons of $\rm{GaAs}$ bilayers recently [34], which allows us to confirm the relevance of the next-neighbor interactions in this Bose–Fermi system.
@@
Figure 3: (a) Phase diagram of the bosonic chemical potential $\mu_{\rm{B}}$ vs. next-neighbor fermion interaction magnitude $V_{\rm{FF}}$ at the thermodynamic limit for a fermionic density $\rho_{\rm{F}}=2/3$ without spin imbalance and local interactions $U_{\rm{FF}}=4$ and $U_{\rm{BF}}=6$. Here we show the evolution of the respective charge gaps from each insulator phase shown in Fig. 2(a) which include the MMI at $\rho_{\rm{B}}=1/3$, the SSMI at $\rho_{\rm{B}}=2/3$ and the ICDW at $\rho_{\rm{B}}=1/2$. The points are extrapolations of the chemical potential to the thermodynamic limit. (b) Order parameter of the ICDW phase $\mathcal{O}^{\rm{ICDW}}$ as a function of $V_{\rm FF}$ in the thermodynamic limit for the same parameters as (a). The vertical dashed line denotes the critical value ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/2}$ of the SF-ICDW transition obtained from the order parameter calculation. The inset shows the dependence of $\mathcal{O}^{\rm{ICDW}}$ on $\rho_{\rm B}$ for a chain of $L=45$ and for $V_{\rm FF}=0$ (black squares) and $V_{\rm FF}=2$ (red circles). Here $\rho_{\rm B}=1/2$ is indicated as a dotted line. Horizontal dashed and dotted lines are added to highlight the zero value of the order parameter. Solid lines are visual guides.
@@
This signals the simultaneous CDW order of both species as the multiplication of the intraspecies CDW order parameters, with a minus sign representing the out-of-phase mutual oscillation, exhibited in Fig. 2(b). Moreover, we subtract the total density from each corresponding density profile to remove the contributions of flat profiles. We show $\mathcal{O}^{\rm{ICDW}}$ for $0\leq V_{\rm FF}\leq 2$ extrapolated to the thermodynamic limit in Fig. 3(b) where we can see the sudden change from zero to positive values after the critical point ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/2}\approx 0.7$ (indicated by the vertical line) which agrees with the one obtained from the gap calculation. The dependence on the number of bosons is also shown in the inset of Fig. 3(b) as we vary $\rho_{\rm B}$ and observe that the main nonzero contribution occurs close to half-filling for $V_{\rm FF}=2$ (red dots) while for $V_{\rm FF}=0$ (black squares) the order parameter is close to zero, as expected. Since we restrict ourselves to odd values of $L$ we observe a deviation of the maximum from $\rho_{\rm B}=1/2$, nevertheless by increasing the system size this finite-size effect disappears as the maximum gets closer to the expected value. Some remarks regarding the order of the transition can be found in Appendix B, a topic which falls outside the scope of this work.
@@
On the other hand, for the MMI at $\rho_{\rm{B}}=1/3$, we see that the charge gap decreases monotonously until it vanishes at a critical value of ${V_{\rm{FF}}^{*}}^{\rho_{\rm{B}}=1/3}\approx 3.5$, as anticipated from the formation of composites that destroy the crystal order mentioned beforehand. For the SSMI, even though we observed an initial increase in its charge gap, here we find that after a certain value, the gap starts to decrease monotonously. This behavior is persistent until the charge gap vanishes at a very high $V_{\rm{FF}}$ value, not shown here. Although the excess of bosonic carriers and the fact that these tie only one kind of fermion together try to maintain a homogeneous distribution of charge in the system, large values of the next-neighbor interaction will force the emergence of compounds that diminish the ground-state energy and finally destroy the crystalline order.
@@
It is worth noting that the ICDW phase does not appear for every fermionic density. Specifically, for $\rho_{\rm{F}}<1/2$, the system does not have enough fermions to build the sub-lattice for the CDW order. This is independent of the spin-population imbalance since the CDW order can be established independently of the number of fermions in each orientation. On the other hand, for $I\neq 0$ there is another limitation. Consider the case where $\rho_{\rm{F}}^{\uparrow}$ or $\rho_{\rm{F}}^{\downarrow}$ are larger than half filling; then all of the fermions with the corresponding spin do not fit in the CDW order, and hence it is not possible to construct the insulating phase. Both restrictions can be summarized with the mathematical condition that the ICDW phase only appears for $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$. The case $I=1$, that is of spin-polarized fermions, has already been reported in the literature [87], and according to this study, there is only one possible density combination in which the insulator can appear, namely $\rho_{\rm{F}}=\rho_{\rm{B}}=1/2$. This emphasizes that a mixture with a larger number of fermionic degrees of freedom admits a more general family of density combinations for the emergence of the ICDW state.
The presence of a non-zero bosonic gap shows the insulating characteristic of this mixture’s component, then it is important to discuss the compressibility of the fermions in the system. Given that $\rho_{\rm B}=1/2$, the conditions for the emergence of the ICDW phase, that is $1/2\leq\rho_{\rm{F}}\leq 1/(1+I)$, draw a region in the $\rho^{\downarrow}_{\rm F}-\rho^{\uparrow}_{\rm F}$ plane, inside which we expect no fermionic charge gap since small deviations to the fermionic densities do not affect the conditions for the insulator to emerge. On the other hand, we do expect a fermionic gap on the boundaries, where the transition to the superfluid phases occurs, as it is known to happen for half-filling ($\rho_{B}=\rho_{F}=1/2$) where the state corresponds to the MMI [68], as an example. This is reminiscent of the supersolid phase where crystalline and superfluid orders are simultaneously present [72, 73], nevertheless a further characterization falls outside the scope of this work.